OpenStudy (luhvon):

what 3/6 written as a decimal?

OpenStudy (superwholock221):

what is 3/6 simplified

OpenStudy (superwholock221):

3 is ____ of 6

OpenStudy (superwholock221):

work with meeeeeeeeee

OpenStudy (mariamikayla):

3/6=1/2=0.5

OpenStudy (superwholock221):

dude your not supposed to give answers

OpenStudy (superwholock221):

luhvon what grade are you in??

OpenStudy (luhvon):

bruh will u please leave them alone like if they give me the answer well oh well let them do what they want to do cuz i can easily inbox them and still get the answer and im in 10th why

OpenStudy (superwholock221):

because this is simple math

OpenStudy (superwholock221):

like im sure if you actually look at it youll understand

OpenStudy (luhvon):

im finna block yo retrice

OpenStudy (superwholock221):

go for it i dont care be childish lol

OpenStudy (luhvon):

are u a girl or a boy cuz if u dude u can shake frl

OpenStudy (superwholock221):

your verrrry scary

OpenStudy (luhvon):

stop talking to me bruh

OpenStudy (superwholock221):

lol no problem "bruh" oh and im a girl

OpenStudy (luhvon):

stop being so harsh towards others that rude u should know

OpenStudy (superwholock221):

im not being harsh im simply saying this site is for help not for cheating and you wanna talk about rude? Threatening other people is rude "if you a dude you can shake frl"

OpenStudy (luhvon):

meaning leave me alone and let others do as they please you cant stop them cuz i still get the answer by inboxing them

OpenStudy (superwholock221):

i was trying to help you understand while others would let you fall on your butt when you dont have a website to cheat with